# Predicting the structure of dynamic graphs

### 08 Jan 2024 v1

@@ 4.1 Synthetic Experiments
Figure 2: Predicted synthetic networks using a sequence of 24 graphs for $h\in\{1,3,5,7\}$ describing graphs at time steps 25, 27, 29 and 31. The predicted graphs have 85, 99, 113, 127 vertices and 470, 567, 643 and 716 edges respectively. Node color depicts the degree of the node, with darker nodes having higher degree.
Figure 2 shows the predicted graphs at $T=24$ for time step $h\in\{1,3,5,7\}$ (predicted graphs at time steps 25, 27, 29 and 31) using $\gamma=0.5$ and $u=0.55$. The number of vertices in the predicted graphs increase with $h$ from 85 to 127, and the predicted edges increase from 470 to 716. Figure 3 shows the adjacency error for graph sequences with $T=20$ for different $h$ values. The results were similar for other $T$ values. We see that as we delete more edges the adjacency error increases, i.e., prediction becomes difficult. This is expected because it is easier to predict repeating patterns than random fluctuations. We also see that adjacency error decreases when $\beta$ (the proportion of edges increased), increases. This is because more dense graphs have more ones in the adjacency matrix, resulting in lower errors.
Figure 3: Results of the synthetic experiments for different time step predictions. The adjacency error of the predicted graphs for $T=20$ for $h\in\{1,\ldots,5\}$ for different values of $\alpha$ (proportions of deleted edges) and $\beta$ (proportion of edges increased).
@@ 4.2 Experiments using real-world datasets
We use the following datasets in this study.
  • UCI Message (UCI): This dataset contains interactions between an online community of students at the University of California Irvine (Panzarasa et al., 2009). Each student is represented by a node and communications between students are represented by edges.
  • EU-email (EU): This dataset was originally investigated by Leskovec et al. (2007) and was later made available by Paranjape et al. (2017). It contains the email network from a European research organisation. An edge $(u,v,t)$ signifies person $u$ sent person $v$ an email at time $t$.
  • High School (HS): This dataset represents contact between students in a high school in Marseilles, France, in December 2013, collected using wearable sensors (Mastrandrea et al., 2015).
  • Hypertext 2009 (HT09): This dataset was collected at the ACM HyperText 2009 Conference, where participants’ face-to-face interactions were recorded using RFID devices (Isella et al., 2011).
Each dataset is an anonymized edgelist of observations of the form $(u,v,t)$. For UCI and EU datasets we consider daily edges and for HS and HT09 datasets we consider hourly edges. Table 1 gives the total number of nodes, edges and density for each dataset for 35 time periods.
Table 1: Dataset summary
Dataset
Total Nodes
Total Edges
Density
UCI
1.3K
25.1K
$7.7\times 10^{-2}$
EU
857
102.6K
$9.7\times 10^{-3}$
HS
327
18.3K
$1.0\times 10^{-2}$
HT09
113
3.6K
$3.3\times 10^{-2}$
Figure 4: UCI-Message predicted networks using a sequence of 20 graphs ($T=20$) for time steps $h\in\{1,3,5\}$. The predicted graphs have 873, 963, 1054 vertices and 924, 1059, 1152 edges respectively. Node color depicts node degree with darker colors depicting higher degree.
Figure 5: Node error of the predicted graphs $\hat{\mathcal{G}}_{T+h}$ for $T\in\{20,\ldots,30\}$ and $h\in\{1,\ldots,5\}$.
Table 2: Results summary
Dataset Form Adjacency Error Degree Cosine Similarity
$h=1$
$h=2$
$h=3$
$h=4$
$h=5$
Err1.
$h=1$
$h=2$
$h=3$
$h=4$
$h=5$
F2C3
1.91 2.03 2.10 2.14 2.15
26.15
0.50
0.44
0.43
0.41
0.38
UCI
F2C2
2.02
2.12
2.18
2.22
2.22
26.15
0.10
0.08
0.05
0.05
0.04
F1C3
3.08
3.35
3.53
3.63
3.69
26.15
0.53 0.48 0.47 0.46 0.43
F1C2
34.40
34.59
34.62
34.33
34.06
26.15
0.33
0.32
0.32
0.33
0.33
F2C3
1.94 1.49 1.56 2.27 2.60
37.28
0.60
0.69 0.67
0.59
0.55
EU
F2C2
2.24
1.83
1.89
2.56
2.87
37.28
0.03
0.03
0.04
0.03
0.02
F1C3
2.67
2.63
2.79
3.63
4.34
37.91
0.62
0.64
0.63
0.60 0.56
F1C2
36.65
32.24
32.32
39.83
45.54
37.91
0.40
0.43
0.43
0.40
0.38
F2C3
1.56 1.54 1.54 1.52 1.46
11.68
0.67
0.67
0.67
0.69
0.69
HS
F2C2
1.83
1.80
1.80
1.78
1.72
11.68
0.16
0.17
0.17
0.18
0.19
F1C3
2.07
2.30
2.48
2.58
2.55
11.68
0.72 0.72 0.70 0.72 0.73
F1C2
11.55
11.57
11.60
11.41
10.82
11.68
0.64
0.65
0.65
0.67
0.68
F2C3
2.98 4.43 4.57 4.69 4.66
20.72
0.61
0.55
0.52
0.50
0.50
HT09
F2C2
3.11
4.54
4.69
4.82
4.79
20.72
0.25
0.22
0.23
0.22
0.16
F1C3
3.87
6.31
7.20
7.71
7.92
20.72
0.68 0.60 0.56 0.53 0.53
F1C2
13.11
19.24
20.03
20.94
20.91
20.72
0.55
0.51
0.49
0.47
0.47
To standardize computations across datasets we predicted graphs $\hat{\mathcal{G}}_{T+h}$ for $T\in\{20,\ldots,30\}$ and $h\in\{1,2,3,4,5\}$. Predictions were carried out for all combinations of $T$ and $h$. We used Formulations 1 and 2 (F1 and F2) with binary and proportional coefficients (C2 and C3). For all configurations we used 0.9 as the upper bound $u$ in $f_{u}(\bm{d})$ and for proportional coefficients C3, we used $\lambda=0.001$.
Figure 4 shows 3 predicted graphs for UCI dataset for $h\in\{1,3,5\}$ at $T=20$ using Formulation 2, coefficients C3, $\gamma=0.5$, $u=0.9$ and $\lambda=0.001$. The nodes and edges in predicted graphs increase with $h$. Figure 5 shows the node error of the datasets for different prediction time-steps $h$. For each time-step $h$ we have 11 predictions for $T\in\{20,\ldots,30\}$ from different graphs $\hat{\mathcal{G}}_{T+h}$. For all datasets apart from High-School dataset we see that the node error increases with the prediction time-step $h$. This is to be expected as it is difficult to forecast further into the future. The High-School dataset is an exception because it has a smaller node growth rate that plateaus after some time.
Table 2 gives the prediction results for these datasets for $\gamma=0.5$ and $u=0.9$. For each dataset and configuration we list the adjacency error and degree cosine similarity for all $h$ values. As there are 11 predictions for each $h$ from different $T$ values we take the average over $T$. As we saw from Table 1 the graph densities are in the $10^{-2}$ range. Thus only this proportion of entries in the adjacency matrix have ones. Therefore, it is quite easy to get an adjacency error greater than 1. To give some perspective of the range of adjacency errors, we compute the error if $\hat{A}_{T+h}-{A}_{T+h}$ is a matrix of all ones. This is not an upper bound because we take the size of $\hat{A}_{T+h}$ to be equal to that of ${A}_{T+h}$ in this computation. However, it is still a useful heuristic because it increases our understanding of the range of adjacency errors. This error is labeled as ERR1 in Table 2.
As general trends we see that the adjacency error increases and the degree cosine similarity decreases as $h$ increases. The configuration F2C3 achieves the best adjacency errors for all 4 datasets followed by F2C2 and F1C3. F1C2 has very large adjacency errors, often greater than ERR1 values. The difference between F1 and F2 is that F2 includes a constraint for total additional edges. The coefficient schemes C2 uses binary coefficients and C3 uses proportional coefficients. In terms of adjacency error, the difference between F2C3 and F1C2 is far greater than changing one setting – either F2 to F1, or C3 to C2 – at a time. The degree cosine similarity is best for F1C3 closely followed by F2C3. We see that F2C2 has poor degree similarity, even though the adjacency error maybe low. From this we can see proportional coefficients (C3) are more important for better predictions than constraining the total edges.
The current formulations solve an optimization problem of $n(n-1)/2$ variables for a graph with $n$ vertices. Thus, the efficiency of the optimization algorithm can be improved. On a different note, we can explore the predictive distribution discussed in Section 3.4 by changing parameters $\gamma$ and $u$. We illustrate graphs with different $\gamma$ and $u$ values in Appendix C.
@@ 5 Conclusion
We have presented a method to predict graphs at future time points from a time series of graphs  [^1] R package netseer available at https://github.com/sevvandi/netseer contains the network prediction algorithm. Programming scripts are available at https://github.com/sevvandi/supplementary_material/tree/master/Graph_Prediction.. Our prediction can include new nodes and edges. The proposed method uses time series methods to predict the node degree and linear programming techniques from flux balance analysis to obtain the future graph. We investigated the utility of our method on synthetic and real datasets. This work focused on predicting the exact graph, i.e., our methodology did not explore assumptions of exchangeability of graph vertices or edges. Investigating prediction methods for exchangeable graphs and improving the efficiency of the optimization algorithm are future research avenues.
Acknowledgments
I sincerely thank Dr Cheng Soon Ong for helpful discussions.
Appendix A Additional methodology details
Theorem 3.2 Let $\mathcal{U}\,,\bm{u}\in\mathcal{U}$ describe the solution space of the optimization problem described in Formulation 1. Then the cardinality of $\mathcal{U}$, denoted by $\left|\mathcal{U}\right|$ satisfies the following bounds
@@ Appendix A Additional methodology details
We consider $\bm{u}_{j}\in\mathcal{U}_{\text{sum}}$ as a binary string. Thus, the number of binary strings in $\mathcal{U}_{\text{sum}}$ is an upper bound for $\left|\mathcal{U}\right|$. From Lemma 3.3 we know $\sum_{k}s_{kl}=\{0,2\}$ for all $k$. As a result we have
@@
Appendix B Synthetic Experimental Results
Figure 6: Degree cosine similarity and node error of synthetic experiments.
Figure 6 shows the degree cosine similarity and node error of synthetic experiments for different time steps $h\in\{1,\ldots,5\}$. As the time step $h$ increases the node error increases and the degree cosine similarity generally decreases. The degree cosine similarity depends on $\alpha$, the proportion of edges deleted and $\beta$ the proportion of edges increased. The degree cosine similarity decreases as $\alpha$ increases but generally increases as $\beta$ increases.
Appendix C Results from Real Datasets
C.1 Graphs from the predictive distribution
Figure 7 shows 3 predicted graphs for UCI-Message dataset for different $\gamma$ and $u$ values. The colors in UCI-Message represent eccentricity of nodes (the maximum distance to other nodes) computed for each node’s connected component with darker colors used for most eccentric nodes and light yellow used for unconnected nodes. The graph predictions are using a time series of graphs with $T=30$ and $h=1$. The leftmost graph shows the predicted network with $\gamma=0.2$ and $u=0.55$. The middle graph correspond to $\gamma=0.5$ and $u=0.95$. The rightmost graph correspond to $\gamma=0.8$ and $u=0.95$. The number of predicted nodes in graphs from left to right are 1113, 1133 and 1166 respectively. This is because the corresponding $\gamma$ values $\{0.2,0.5,0.8\}$ give rise to different number of nodes in the predicted graph. The predicted graphs from left to right have 494, 646, and 646 edges respectively. This is a result of the $u$ values $\{0.55,0.95,0.95\}$. Even though $\gamma$ values in the middle and the rightmost graph are different (0.5 and 0.8) the number of edges is the same because we are using Formulation 2, which constraints the total number of edges to the predicted total number of edges. Had we used Formulation 1, the edges would have increased as well. The illustrated graphs are specific instances of the distribution of predicted graphs, and similar graphs can be obtained by varying $\gamma$ and $u$ values.
Figure 7: UCI-Message predicted networks for different values of $\gamma$ and $u$. Nodes colored by node eccentricity with yellow for unconnected and blue for most eccentric nodes.
Figure 8: EU-Email predicted networks for $\gamma=0.5$ and $u\in\{0.55,0.80,0.95\}$. Nodes coloured by node eccentricity.
Figure 8 shows 3 predicted graphs for the EU-Email dataset for $\gamma=0.5$ and $u\in\{0.55,0.80,0.95\}$ from a time series of graphs with $T=30$ and $h=1$. Again nodes are colored by their eccentricity with unconnected nodes colored the lightest and most eccentric nodes colored the darkest. All 3 predicted graphs have 706 nodes corresponding to $\gamma=0.5$. The number of edges in the leftmost, middle and the rightmost graphs are 916, 1118 and 1377 respectively. The number of edges increase because the upper bound $u$ increased from left to right.

### 25 Jul 2024 v2

@@ 2 Node C5 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
Figure 3: Forecast synthetic networks using a sequence of 25 networks for $h\in\{1,3,5\}$ illustrating the forecast graphs at time steps 26, 28 and 30. Node colour depicts the degree of the node, with darker nodes having a higher degree.
Table 1 gives the evaluation metrics for the first and second experiments. We see that for both experiments coefficient schemes C5 and C6 achieve better results compared to the last seen network, which is denoted by LS. The node error of C5 and C6 coefficient schemes is 0 for both experiments. The edge error is quite small for both C5 and C6 compared to LS. Similarly the edge density error for C5 and C6 is much smaller than that of LS.
Figure 3 illustrates the graph forecasts for synthetic experiment 2 for time steps $h\in\{1,3,5\}$. These are graph forecasts at time steps 26, 28 and 30 using $\gamma=0.5$ and $u=0.55$. The node color reflects the degree of the node with darker shades denoting nodes with higher degree. We see the algorithm has accounted for graph growth.
@@ 4.2 Experiments using real-world datasets
We use two realworld datasets in this study:
  1. 1.
    Facebook (FB): This dataset contains users and their links from the Facebook New Orleans networks (Viswanath et al., 2009). Users are depicted as nodes in the network and an edge exists between two users if they are friends. There are timestamped and non-timestamped edges in this dataset. We used timestamped edges for network forecasting.
  2. 2.
    Physics citations (Hep-PH): This dataset contains citations of high energy physics papers published on ArXiv (Leskovec et al., 2007). Each paper is denoted as a node and citations between papers are denoted as edges.
Both FB and Hep-PH datasets are anonymized edge lists of observations of the form $(u,v,t)$. For FB we consider daily edges to form a network and for Hep-PH we consider monthly edges. For both networks $E_{t}\not\subset E_{t+1}$ because only edges in each time window are used to construct the network. Figure 4 shows the number of edges and vertices in each network $\mathcal{G}_{t}$ for $t\in\{1,\ldots,35\}$.
To standardize computations across the two datasets we consider the network time series $\{\mathcal{G}_{t}\}_{t=1}^{T}$ with $T\in\{20,\ldots,30\}$ and forecast $\hat{\mathcal{G}}_{T+h}$ for $h\in\{1,\ldots,5\}$. Forecasts were carried out for all combinations of $T$ and $h$. We used Formulation 2 (F2) with harmonic decay (C5) and last seen graph coefficients (C6) to forecast the network. For all configurations we used 0.55 as the upper bound $u$ in $f_{u}(\bm{d})$. As a means of comparison, we used the last seen (LS) graph $\mathcal{G}_{T}$ and computed the evaluation metrics for LS.
Table 2 shows the graph forecasting results for FB and Hep-PH datasets. The node, edge and density error for coefficient schemes C5 and C6 are given along with the metrics of the last seen (LS) graph. We see that for almost all timesteps $h$, C5 and C6 perform better than LS in all metrics. The only exception is the edge error for $h=1$ in Hep-PH, for which LS achieves a lower value than C5 or C6. For both FB and Hep-PH, the node error is much lower for C5 and C6 compared to LS. The edge error too is lower for C5 and C6. All edge density errors are low as these graphs are sparse even though C5 and C6 achieve lower errors. We also see that several timesteps ahead forecasts have larger errors compared to forecasts with shorter timesteps, which is to be expected.
Figure 4: Edges and vertices in each network for realworld datasets.
Table 2: Results for FB and Hep-PH datasets. Coefficient schemes C5 and C6 are considered. The last seen graph (LS) is considered as a means of comparison. Evaluation metrics node, edge and density errors are given with mean and standard deviation in brackets for 10 iterations.
Dataset Metric Coef. $h=1$ $h=2$ $h=3$ $h=4$ $h=5$
FB Node C5 0.0064(0.0062) 0.0126(0.0115) 0.0195(0.0155) 0.0275(0.0176) 0.0315(0.0194)
Err. C6 0.0064(0.0062) 0.0126(0.0115) 0.0195(0.0155) 0.0275(0.0176) 0.0315(0.0194)
LS 0.0329(0.0076) 0.0641(0.0128) 0.0917(0.0188) 0.1162(0.0244) 0.1392(0.0271)
Edge C5 0.1379(0.0797) 0.1514(0.1020) 0.1826(0.1143) 0.1686(0.1156) 0.1763(0.1177)
Err. C6 0.1353(0.0777) 0.1530(0.1045) 0.1826(0.1143) 0.1686(0.1156) 0.1763(0.1177)
LS 0.1586(0.1047) 0.2348(0.1689) 0.3097(0.1805) 0.2812(0.1750) 0.2107(0.2049)
Dens C5 4e-06(2e-06) 5e-06(3e-06) 5e-06(3e-06) 5e-06(3e-06) 6e-06(5e-06)
Err. C6 4e-06(1e-06) 5e-06(3e-06) 5e-06(3e-06) 5e-06(3e-06) 6e-06(5e-06)
LS 5e-06(3e-06) 8e-06(5e-06) 1e-05(5e-06) 9e-06(6e-06) 8e-06(6e-06)
Hep-PH Node C5 0.0120(0.0083) 0.0157(0.0156) 0.0237(0.0188) 0.0325(0.0258) 0.0441(0.0319)
Err. C6 0.0120(0.0083) 0.0157(0.0156) 0.0237(0.0188) 0.0325(0.0258) 0.0441(0.0319)
LS 0.0775(0.0148) 0.1458(0.0241) 0.2064(0.0329) 0.2594(0.0406) 0.3084(0.0473)
Edge C5 0.2250(0.1374) 0.2139(0.2028) 0.2314(0.1192) 0.1588(0.1484) 0.2380(0.1346)
Err C6 0.2250(0.1374) 0.2139(0.2028) 0.2314(0.1192) 0.1588(0.1484) 0.2380(0.1346)
LS 0.1973(0.1174) 0.2364(0.1665) 0.2489(0.1508) 0.2361(0.1737) 0.3113(0.2063)
Dens. C5 4e-05(3e-05) 3e-05(3e-05) 3e-05(2e-05) 2e-05(2e-05) 3e-05(2e-05)
Err. C6 4e-05(3e-05) 3e-05(3e-05) 3e-05(2e-05) 2e-05(2e-05) 3e-05(2e-05)
LS 4e-05(3e-05) 5e-05(4e-05) 5e-05(4e-05) 6e-05(4e-05) 6e-05(4e-05)
@@ 5 Conclusion
We have presented a method to forecast graphs at future time points from a time series of graphs. Our forecast can include new nodes and edges. The proposed method uses time series methods to forecast the node degree and linear programming techniques from flux balance analysis to obtain future graphs. The R package netseer(Sevvandi Kandanaarachchi, n.d.) includes the methodology. We investigated the utility of our method on synthetic and real datasets. This work focused on unweighted graphs. Future research avenues include extending this methodology to weighted and directed graphs.
Appendix: Proof of Theorem 3.2
@@
We consider $\bm{u}_{j}\in\mathcal{U}_{\text{sum}}$ as a binary string. Thus, the number of binary strings in $\mathcal{U}_{\text{sum}}$ is an upper bound for $\left|\mathcal{U}\right|$. As $S$ is a zero-inflated incidence matrix, $\sum_{k}s_{kl}=\{0,2\}$ for all $k$. As a result we have